Journalist Report – January 3rd

MDRS Crew 202, Journalist Report

Sol 5 – 01/03/2018

Name the space movie (or show) given the following quote. Answer at the end of the Report:

So, yes, they let women do some things at NASA, Mr. Johnson, and it’s not because we wear skirts. It’s because we wear glasses. Have a good day.

Update on the Great Generator Crisis of Crew 202. It was not in fact the generator, but the power system itself. The main controller’s personality took on a Hal 2000 fashion, trying to freeze us inside of our own home. I am happy to report that our wonderful MDRS Director and Assistant Director have diagnosed the issue and we are no longer in low power mode. Today, we can use the oven again.

Day to day life on Mars is a semi-structured routine. You wake up at 7:30 am to a crew selected song. This morning was “Pina Coladas”, a drink which the crew enjoys, none of which is on Mars. At 8:30 am, we start a yoga/stretching/dynamic exercise session, traditionally led by our Engineer, who is a rowing champion back in her former role at Purdue, and our Executive Officer, who, after the mission, will return to Earth to resume his role as an Army drill sergeant. Once our blood is pumping, breakfast is started by two crew members while the others plan and prepare for the day’s activities. This morning’s breakfast was led by our Geologist and assisted by myself. With a box of oats, dehydrated apples, a hefty spoonful of brown sugar, and a dash of cinnamon, the crew enjoyed a bowl of apple cinnamon oatmeal that rivaled even the best Quaker Oat packets (obviously the dinosaur eggs flavor). The majority of our agendas will involve an Extra-Vehicular Activity (EVA) with 3-4 crew members. Suit up at 10 am. In the airlock by 11 am, and away the crew goes to make a dent into their Martian research.

The Engineer, the Executive Officer, and myself traveled to an area commonly called “The Kissing Camels”. It is these huge hills with brilliant stripes of oranges, reds, and yellows whose sides turn into cliffs and extend in a semi-circle around a plain. Our mission today: Collect various rocks and boulders for our Geologist to analyze, gather radioactive data across the area, and spit in a cup upon our return for a stress study. On a future Mars mission, we anticipate the crew would be trained to identify different rock formations as geological studies are of great interest in Martian research. This is not a future mission and no one on the EVA is a geologist. We had a briefing with our Geologist and our Commander, who was the Geologist on a previous mission, on what to look for in the rocks, all of the different textures, colors, and strengths. They were interested in “fine coarse” shale, sand found near former riverbeds, and different colored boulders, specifically red/brown and green/blue. These objectives seemed easy enough until you start questioning your very understanding of colors. Is this rock truly red? Is this the green our Geologist was talking about? What even is Blue? At one point, we simply switched to the method “This rock looks interesting”, threw it in a Ziploc sampling bag, and continued on our way. Thankfully, our Geologist appeared pleased with our haul.

The Health and Safety Officer (HSO), Geologist, and Commander spent the day making progress on individual research projects. The HSO has been working like a machine in the green habitat. He’s busy prepping his 10 day germination and growth research, sterilizing dishes and hydroponically planting his microgreens. Hydroponics means no soil which is ideal for a Martian environment where the soil may not be suitable for growing and it is a long way to ship from Earth! Our Geologist is continuing a MDRS study from 2014 and is preparing for an EVA tomorrow to collect additional geological samples. Our Commander is commanding, keeping the ship running, and ensuring everyone has what they need, to do what they need to do. With the success of the power being restored and the progress being made on our ambitious research goals, the crew morale is high on our sixth mission day.

Tonight, we settled into a gumbo dinner and plan to indulge in a pineapple upside down cake made from our Mars friendly ingredients. Baking is strange when everything needs to be re-hydrated. It creates an odd spongy texture, but the flavor is still delicious and sweet.

Here at MDRS, we are far from civilization. The closest town is only populated by ~200 people and 30 minutes down an off-road trail through plains and mountains. In our Great Generator Crisis, the low power mode forced us to turn off all unnecessary electricity. The Green Habitat, the Science Dome, and the majority of lights in the Habitat were all shut down to conserve power. This brief inconvenience turned into a beautiful accident by the moonless sky that appeared above us. I’ve never seen so many stars in all their varying brightness and depths. Even the milky way was unveiled, as if someone took a paint brush of white paint from one side of the horizon, across the highest point in the sky, to the other behind the mountains in the distance. It. Was. Breathtaking. It makes us wonder, what night sky will future Martians see from their base? Will they gaze at Earth with as much wonderment and excitement as we view Mars, shining a subdued red in the night? When we gaze at the constellations, viewable from both planets, will future Martians feel comforted or homesick?

Movie (or Show) Answer: Hidden Figures

Astronomy Report – January 3rd

Crew 202 Astronomy Report 03-Jan-2019

Crew Astronomer: Cesare Guariniello

MDRS ROBOTIC OBSERVATORY

Robotic Telescope: MDRS-WF

Objects Viewed: M51, M81 and M82, M31

Problems Encountered: none (photos will be sent after post-processing)

MUSK OBSERVATORY

Solar Features Observed: Three sunspots (two in a group), optical observation only

Problems Encountered: The first centering needed quite a large amount of manual adjustment.

Operations Report – January 3rd

Crew 202 Operations Report 02-JAN-2019

SOL: 5

Name of person filing report: Kasey Hilton

Non-nominal systems: Spacesuit charger

Notes on non-nominal systems: One of the Tenergy brand spacesuit chargers stopped working. It looked like the inside of the charger plug had been burnt. It was replaced with a PeleusTech charger. There are no more back up chargers in the suit supply box in the RAM.

Generator (hours run): 1hr 34min; Turned on last night (02Jan2019) at 17:45 until 19:19 when the power system program malfunctioned. The generator did not run overnight.

Solar – SOC 83%

Diesel Reading – 73%

Propane Reading – 50%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 75%, about a foot and a half below the top; 430 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 80%; 240 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 10

Static to Loft Pump used – Yes; At 10:00 to refill the tank

Water Meter: 01397842 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 64.2

Beginning charge: 100%

Ending charge: 68%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 62.4

Beginning charge: 100%

Ending charge: 50%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: A power strip added to the EVA room to make all the chargers reach the suits.

Summary of internet: Nothing to report

Summary of suits and radios: One of the suit chargers stopped working and was replaced with a backup charger that was in the RAM as mentioned above in “Non-nominal systems”. The replacement charger was the last backup that was in the spacesuit supply box.

Summary of Hab operations: Hab is out of power conservation mode. Power is now nominal.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Greenhab Report – January 3rd

Crew 202 Greenhab Report 03-Jan-2019
Greenhab Officer:
Jake Qiu

Environmental Control:
Heating
Cooling w/ ambient air (1hr)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 22.9°C
High: 40.0°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
16 gallons

Water in Blue Tank (lbs):
86%

Times of Watering for Crops:
1014
1855

Changes to crops:
* 7 growth trays of daikon radish micro-greens prepared
* Sunflowers Growing very well
* Some of savory mix greens failing to sprout
* Radishes growing very well
* Tomatoes still lacking growth

Narrative:
Finished preparing the 7 growth trays for control condition. Placed all of the plants in its original location.
Placed the 80% shade back on in its original location as well. Watered all of the plants as needed and some of
the plants are showing good progress.

Experiment Results:
Date: SOL5 Crew 202 (SOL14 overall)

Prepared 7 growth trays as control condition. Sampled the micro-greens for grey water and soil conditions.
Watered all 21 trays as needed.

Watered previous crew 201 experiments

Harvest: N/A

Support/Supplies Needed: N/A

Science Report – January 3rd

Science Report
3 Jan 2019
Crew 202 – MartianMakers
Submitted by GEO Ellen Czaplinski

Analysis of mineralogy and regime of sand dunes and fluvial processes: The crew collected geological samples from six sites during their EVA near Kissing Camel Range. Samples included whole rocks, rocks separated from a larger parent body, and sand samples. Initial sample analysis consisted of classifying the rocks based on texture, grain size, sorting, and color using a hand lens and microscope. The samples were primarily sandstone, clays, shale, and sand from fluvial features like channels. Further analysis will need to be carried out in order to determine the mineralogy of the samples with the TREK portable spectrometer, and eventually compared to IR spectra of the samples taken in the lab.

EVA Report 4 – January 3rd

Crew 202 EVA Report 3-Jan-2019

EVA #4

Author: Denys Bulikhov (EXO)

Purpose of EVA: Collection of geological samples and ambient radiation readings

Start time: 10:50

End time: 12:44

Narrative: EVA 4 crew took Cow Dung Road down to the Kissing Camels Ridge. Crew went to the west side of the Ridge first collecting samples from fallen rocks and river beds and collecting ambient radiation readings. Then they went to the East side of the Ridge collecting samples and readings, made a circle around the valley and came back to the rovers.

Destination: Kissing Camels Ridge

Coordinates: N518500, E4249300

Participants: Denys Bulikhov (EXO), Kasey Hilton (ENG), Alexandra Dukes (JOU)

Road(s) and routes per MDRS Map: Cow Dung Road, then walk along the Kissing Camels Ridge.

Mode of travel: Driving and walking

Vehicles used: Spirit and Curiosity.

Sol Summary 5 – January 3rd

Crew 202 Sol Summary Report 03-Jan-2019

Sol: 5

Summary Title: After a low-power night

Author’s name: Cesare Guariniello

Mission Status: The crew is continuing work on all research project. The morning was spent in power conservation mode, during which half of the crew performed an EVA and the other half stayed in the habitat to work on various projects. The power is now nominal.

Sol Activity Summary: After a good morning stretching and some light exercises, half of the crew went on EVA #4 while the rest worked on projects (and some origami) in the habitat. The afternoon was spent resting and working on research projects, including microbiology, geology, and astronomy. The crew is doing great, and every little difficulty or tiredness is overcome with laughter and a cohesion among crew members which is a joy to the eyes!

Look Ahead Plan: Tomorrow we will have a geology EVA in areas of sand dunes and continue our projects.

Anomalies in work: None

Weather: Sunny and cold with melting snow on the ground

Crew Physical Status: In perfect health

EVA: The Executive Officer, Crew Engineer, and Crew Journalist had an EVA to the region North of Kissing Camel Range, to collect clay and shale samples and identify sandy regions.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, EVA report, EVA request, Geology Report, Journalist Report

Support Requested: None

Cesare
Commander, MDRS 202